OpenStudy (ledah):

suppose y varies jointly with the cube of A and the square of B. When A=4 and B=3,y=10,368. Find y when A=5 and B=4 A:36,000 B:3600 C:18 D:17,280

OpenStudy (radar):

\[y=kA ^{3}B ^{2}\]Where k is the constant of proportionality. \[10,368=k4 ^{3}3^{2}\]\[10,368=k(64)(9)\]k=10368/576=18. This gave us the value of k. Now we use k to solve the rest of the problem y=18(5^3)(4^2)=18*125*16=36,000